\begin{thebibliography}{10}

\bibitem{AGLERMCCARTHY}
Jim Agler and John~E. McCarthy.
\newblock {\em Pick interpolation and {H}ilbert function spaces}, volume~44 of
  {\em Graduate Studies in Mathematics}.
\newblock American Mathematical Society, Providence, RI, 2002.

\bibitem{AKCOGLUSUCHESTON}
M.~A. Akcoglu and L.~Sucheston.
\newblock Dilations of positive contractions on {$L_{p}$} spaces.
\newblock {\em Canad. Math. Bull.}, 20(3):285--292, 1977.

\bibitem{AKCOGLUKOPP}
Mustafa~A. Akcoglu and P.~Ekkehard Kopp.
\newblock Construction of dilations of positive {$L_{p}$}-contractions.
\newblock {\em Math. Z.}, 155(2):119--127, 1977.

\bibitem{AMBROZIEAMULLER}
C.~Ambroziea and V.~Muller.
\newblock Commutative dilation theory.
\newblock In {\em Operator Theory}, pages 1--29. Springer, 2015.

\bibitem{ANDO}
T.~And\^{o}.
\newblock On a pair of commutative contractions.
\newblock {\em Acta Sci. Math. (Szeged)}, 24:88--90, 1963.

\bibitem{ARVESON}
William Arveson.
\newblock Dilation theory yesterday and today.
\newblock In {\em A glimpse at {H}ilbert space operators}, volume 207 of {\em
  Oper. Theory Adv. Appl.}, pages 99--123. Birkh\"{a}user Verlag, Basel, 2010.

\bibitem{BERCOVICI}
Hari Bercovici.
\newblock {\em Operator theory and arithmetic in {$H^\infty$}}, volume~26 of
  {\em Mathematical Surveys and Monographs}.
\newblock American Mathematical Society, Providence, RI, 1988.

\bibitem{BHATDERAKSHITH}
B.~V.~Rajarama Bhat, Sandipan De, and Narayan Rakshit.
\newblock A caricature of dilation theory.
\newblock {\em Adv. Oper. Theory}, 6(4):Paper No. 63, 20, 2021.

\bibitem{BHATMUKHERJEE}
B.~V.~Rajarama Bhat and Mithun Mukherjee.
\newblock Two states.
\newblock {\em Houston J. Math.}, 47(1):63--95, 2021.

\bibitem{BHATTACHARYYA}
Tirthankar Bhattacharyya.
\newblock Dilation of contractive tuples: a survey.
\newblock In {\em Surveys in analysis and operator theory ({C}anberra, 2001)},
  volume~40 of {\em Proc. Centre Math. Appl. Austral. Nat. Univ.}, pages
  89--126. Austral. Nat. Univ., Canberra, 2002.

\bibitem{CHOIDAVIDSON}
Man-Duen Choi and Kenneth~R. Davidson.
\newblock A {$3\times 3$} dilation counterexample.
\newblock {\em Bull. Lond. Math. Soc.}, 45(3):511--519, 2013.

\bibitem{CRABBDAVIE}
M.~J. Crabb and A.~M. Davie.
\newblock von {N}eumann's inequality for {H}ilbert space operators.
\newblock {\em Bull. London Math. Soc.}, 7:49--50, 1975.

\bibitem{DOUGLAS}
R.~G. Douglas.
\newblock Structure theory for operators. {I}.
\newblock {\em J. Reine Angew. Math.}, 232:180--193, 1968.

\bibitem{DRURY}
S.~W. Drury.
\newblock Remarks on von {N}eumann's inequality.
\newblock In {\em Banach spaces, harmonic analysis, and probability theory
  ({S}torrs, {C}onn., 1980/1981)}, volume 995 of {\em Lecture Notes in Math.},
  pages 14--32. Springer, Berlin, 1983.

\bibitem{DURSZTNAGY}
E.~Durszt and B.~Sz.-Nagy.
\newblock Remark to a paper: ``{M}odels for noncommuting operators'' [{J}.
  {F}unct. {A}nal. {\bf 48} (1982), no. 1, 1--11] by {A}. {E}. {F}razho.
\newblock {\em J. Functional Analysis}, 52(1):146--147, 1983.

\bibitem{EGERVARY}
E.~Egerv\'{a}ry.
\newblock On the contractive linear transformations of {$n$}-dimensional vector
  space.
\newblock {\em Acta Sci. Math. (Szeged)}, 15:178--182, 1954.

\bibitem{FACKLERGLUCK}
Stephan Fackler and Jochen Gluck.
\newblock A toolkit for constructing dilations on {B}anach spaces.
\newblock {\em Proc. Lond. Math. Soc. (3)}, 118(2):416--440, 2019.

\bibitem{FENDLER}
Gero Fendler.
\newblock On dilations and transference for continuous one-parameter semigroups
  of positive contractions on {${L}^p$}-spaces.
\newblock {\em Ann. Univ. Sarav. Ser. Math.}, 9(1):iv+97, 1998.

\bibitem{FOIASFRAZHOGOHBERGKAASHOEK}
C.~Foias, A.~E. Frazho, I.~Gohberg, and M.~A. Kaashoek.
\newblock {\em Metric constrained interpolation, commutant lifting and
  systems}, volume 100 of {\em Operator Theory: Advances and Applications}.
\newblock Birkh\"{a}user Verlag, Basel, 1998.

\bibitem{FOIASFRAZHO}
Ciprian Foias and Arthur~E. Frazho.
\newblock {\em The commutant lifting approach to interpolation problems},
  volume~44 of {\em Operator Theory: Advances and Applications}.
\newblock Birkh\"{a}user Verlag, Basel, 1990.

\bibitem{FRAZHO}
Arthur~E. Frazho.
\newblock Models for noncommuting operators.
\newblock {\em J. Functional Analysis}, 48(1):1--11, 1982.

\bibitem{HALMOS}
Paul~R. Halmos.
\newblock Normal dilations and extensions of operators.
\newblock {\em Summa Brasil. Math.}, 2:125--134, 1950.

\bibitem{HANLARSONLIULIU}
Deguang Han, David~R. Larson, Bei Liu, and Rui Liu.
\newblock Structural properties of homomorphism dilation systems.
\newblock {\em Chin. Ann. Math. Ser. B}, 41(4):585--600, 2020.

\bibitem{KERNNAGELPALM}
Martin Kern, Rainer Nagel, and Gunther Palm.
\newblock Dilations of positive operators: construction and ergodic theory.
\newblock {\em Math. Z.}, 156(3):265--277, 1977.

\bibitem{MAHESH}
K.~Mahesh Krishna.
\newblock p-adic magic contractions, p-adic von {N}eumann inequality and p-adic
  {S}z.-{N}agy dilation.
\newblock {\em arXiv:2209.12012v1 [math.NT] 24 September}, 2022.

\bibitem{KRISHNAJOHNSON}
K.~Mahesh Krishna and P.~Sam Johnson.
\newblock Dilations of linear maps on vector spaces.
\newblock {\em Oper. Matrices}, 16(2):465--477, 2022.

\bibitem{LEVYSHALIT}
Eliahu Levy and Orr~Moshe Shalit.
\newblock Dilation theory in finite dimensions: the possible, the impossible
  and the unknown.
\newblock {\em Rocky Mountain J. Math.}, 44(1):203--221, 2014.

\bibitem{MCCARTHYSHALIT}
John~E. McCarthy and Orr~Moshe Shalit.
\newblock Unitary {$N$}-dilations for tuples of commuting matrices.
\newblock {\em Proc. Amer. Math. Soc.}, 141(2):563--571, 2013.

\bibitem{MILNORHUSEMOLLER}
John Milnor and Dale Husemoller.
\newblock {\em Symmetric bilinear forms}.
\newblock Ergebnisse der Mathematik und ihrer Grenzgebiete, Band 73.
  Springer-Verlag, New York-Heidelberg, 1973.

\bibitem{NAGELPALM}
Rainer Nagel and G\"{u}nther Palm.
\newblock Lattice dilations of positive contractions on {$L^{p}$}-spaces.
\newblock {\em Canad. Math. Bull.}, 25(3):371--374, 1982.

\bibitem{PARROTT}
Stephen Parrott.
\newblock Unitary dilations for commuting contractions.
\newblock {\em Pacific J. Math.}, 34:481--490, 1970.

\bibitem{PAULSENBOOK}
Vern Paulsen.
\newblock {\em Completely bounded maps and operator algebras}, volume~78 of
  {\em Cambridge Studies in Advanced Mathematics}.
\newblock Cambridge University Press, Cambridge, 2002.

\bibitem{PISIERBOOK}
Gilles Pisier.
\newblock {\em Similarity problems and completely bounded maps}, volume 1618 of
  {\em Lecture Notes in Mathematics}.
\newblock Springer-Verlag, Berlin, 2001.

\bibitem{POPESCU}
Gelu Popescu.
\newblock Isometric dilations for infinite sequences of noncommuting operators.
\newblock {\em Trans. Amer. Math. Soc.}, 316(2):523--536, 1989.

\bibitem{SARASON}
Donald Sarason.
\newblock Generalized interpolation in {$H^{\infty }$}.
\newblock {\em Trans. Amer. Math. Soc.}, 127:179--203, 1967.

\bibitem{SCHAFFER}
J.~J. Sch\"{a}ffer.
\newblock On unitary dilations of contractions.
\newblock {\em Proc. Amer. Math. Soc.}, 6:322, 1955.

\bibitem{ORRGUIDED}
Orr~Moshe Shalit.
\newblock Dilation theory: a guided tour.
\newblock In {\em Operator theory, functional analysis and applications},
  volume 282 of {\em Oper. Theory Adv. Appl.}, pages 551--623.
  Birkh\"{a}user/Springer, Cham, 2021.

\bibitem{STROESCU}
Elena Stroescu.
\newblock Isometric dilations of contractions on {B}anach spaces.
\newblock {\em Pacific J. Math.}, 47:257--262, 1973.

\bibitem{NAGY}
B\'{e}la Sz.-Nagy.
\newblock Sur les contractions de l'espace de {H}ilbert.
\newblock {\em Acta Sci. Math. (Szeged)}, 15:87--92, 1953.

\bibitem{NAGY1960}
B\'{e}la Sz.-Nagy.
\newblock On {S}ch\"{a}ffer's construction of unitary dilations.
\newblock {\em Ann. Univ. Sci. Budapest. E\"{o}tv\"{o}s Sect. Math.},
  3(4):343--346, 1960/61.

\bibitem{NAGYLIFTING}
B\'{e}la Sz.-Nagy and Ciprian Foia\c{s}.
\newblock The ``lifting theorem'' for intertwining operators and some new
  applications.
\newblock {\em Indiana Univ. Math. J.}, 20(10):901--904, 1971.

\bibitem{NAGYFOIAS}
Bela Sz.-Nagy, Ciprian Foias, Hari Bercovici, and Laszlo Kerchy.
\newblock {\em Harmonic analysis of operators on {H}ilbert space}.
\newblock Universitext. Springer, New York, second edition, 2010.

\bibitem{VAROPOULOS}
N.~Th. Varopoulos.
\newblock On an inequality of von {N}eumann and an application of the metric
  theory of tensor products to operators theory.
\newblock {\em J. Functional Analysis}, 16:83--100, 1974.

\end{thebibliography}

\documentclass{amsart}
\usepackage{setspace}
\usepackage{a4}
\usepackage{amssymb,amsmath,amsthm,latexsym}
\usepackage{amsfonts}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{cite}
\usepackage{enumerate}
\usepackage[mathscr]{euscript}
\usepackage{mathtools}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{definition}[theorem]{Definition}

\newtheorem{problem}[theorem]{Problem}

\title{This is the title}
\usepackage{amssymb}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{hyperref}
\usepackage{enumerate}
\usepackage{mathtools}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{hyperref}
\usepackage{amssymb}
\usepackage{enumerate}
\usepackage{mathtools}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{graphicx}
\usepackage{textcomp}
\usetikzlibrary{cd}
\usepackage{fancyhdr}

\begin{document}
	\hrule\hrule\hrule\hrule\hrule
	\vspace{0.3cm}	
	\begin{center}
		{\bf{INDEFINITE HALMOS, EGERVARY AND   SZ.-NAGY  DILATIONS}}\\
		\vspace{0.3cm}
		\hrule\hrule\hrule\hrule\hrule
		\vspace{0.3cm}
		\textbf{K. MAHESH KRISHNA}\\
		Post Doctoral Fellow \\
		Statistics and Mathematics Unit\\
		Indian Statistical Institute, Bangalore Centre\\
		Karnataka 560 059, India\\
		Email: kmaheshak@gmail.com\\
		
		Date: \today
	\end{center}

\hrule
\vspace{0.5cm}
%--------------------------------------
\textbf{Abstract}: Let $\mathcal{M}$ be an indefinite inner product module over a *-ring of characteristic 2. We show that every self-adjoint operator on $\mathcal{M}$ admits Halmos, Egervary and Sz.-Nagy dilations.

\textbf{Keywords}:  Dilation, Indefinite inner product space, Module.

\textbf{Mathematics Subject Classification (2020)}: 47A20, 16D10, 46C20.
\vspace{0.5cm}
\hrule 
%\tableofcontents

\section{Introduction}
In 1950,  Halmos \cite{HALMOS}  made a deep insight into structure theory of operators on Hilbert space by exhibiting any contraction as a part of a unitary. In 1953, Sz.-Nagy \cite{NAGY} showed that Halmos result can be extended to powers of contractions using a unitary operator. In 1963, T. Ando \cite{ANDO} showed that there is a version of Sz.-Nagy dilation for commuting contractions. Combined with spectral theory and theory of (several) complex variables, today, dilation theory of contractions is a rapidly evolving area of research and for a comprehensive look, we refer  \cite{SCHAFFER, NAGYLIFTING, ANDO, PAULSENBOOK, PISIERBOOK, SARASON, AMBROZIEAMULLER, AGLERMCCARTHY, DOUGLAS, NAGYFOIAS, LEVYSHALIT, ARVESON, ORRGUIDED, FRAZHO, FOIASFRAZHO, FOIASFRAZHOGOHBERGKAASHOEK, DURSZTNAGY, POPESCU, BERCOVICI, BHATMUKHERJEE, EGERVARY, BHATTACHARYYA, NAGY1960, PARROTT, DRURY, CRABBDAVIE, MCCARTHYSHALIT, VAROPOULOS, CHOIDAVIDSON}. Started in 1970's, dilations of contractions acting on Lebesgue spaces and  Banach spaces  followed Hilbert space developments \cite{FACKLERGLUCK, AKCOGLUSUCHESTON, STROESCU, NAGELPALM, AKCOGLUKOPP, KERNNAGELPALM, FENDLER}.

In 2021, by identifying essential mechanisms of dilation theory, Bhat,  De and Rakshit \cite{BHATDERAKSHITH} obtained surprising results in the set theory context and vector spaces. In 2022, further study in the context of vector spaces was carried  by Krishna and Johnson \cite{KRISHNAJOHNSON}. We note that another vector space variant is also studied by 
Han, Larson, Liu and Liu \cite{HANLARSONLIULIU}. Recently Krishna introduced the notion of magic contractions and derived Sz.-Nagy dilation for p-adic Hilbert spaces and modules \cite{MAHESH}.

In this paper, we  derive indefinite inner product module  versions of Halmos dilation (Theorem \ref{HD}), Egervary N-dilation (Theorem \ref{ED}), Sz.-Nagy dilation (Theorem \ref{ND}). Our article is highly motivated from the paper of Halmos \cite{HALMOS}, Egervary \cite{EGERVARY}, Schaffer \cite{SCHAFFER}, Sz.-Nagy \cite{NAGY}, Bhat, De and Rakshit \cite{BHATDERAKSHITH}, Krishna and Johnson \cite{KRISHNAJOHNSON} and   Krishna \cite{MAHESH}.

\section{Indefinite Halmos, Egervary and Sz.-Nagy  Dilations}
We are going to use the following notions. A ring $\mathcal{R}$ with an automorphism $*$ which is either identity or of order 2 is called as an *-ring. Throughout the paper we assume that characteristic of ring is 2. 
\begin{definition}\cite{MILNORHUSEMOLLER}\label{PADICDEF}
Let  $\mathcal{V}$ be a module  over $\mathcal{R}$. We say that $\mathcal{V}$ is an indefinite inner product module  (we write IIPM) if there is a map (called as indefinite inner product) $\langle \cdot, \cdot \rangle: \mathcal{V} \times \mathcal{V} \to \mathcal{R}$ satisfying following.
\begin{enumerate}[\upshape (i)]
	\item If $x \in \mathcal{V}$ is such that $\langle x,y \rangle =0$ for all $y \in \mathcal{V}$, then $x=0$.
	\item $\langle x, y \rangle =\langle y, x \rangle^*$ for all $x,y \in \mathcal{V}$.
	\item $\langle \alpha x+y, z \rangle =a \langle x,  z \rangle +\langle y,z \rangle $ for all  $a  \in \mathcal{R}$, for all $x, y, z \in \mathcal{V}$.
\end{enumerate}
\end{definition}
Let $\mathcal{V}$ be a  IIPM and $T:\mathcal{V}\to \mathcal{V}$ be a morphism. We say that $T$ is adjointable if there is a morphism, denoted by $T^*:\mathcal{V}\to \mathcal{V}$ such that $\langle Tx,y\rangle =\langle x,T^*y\rangle$, $\forall x, y \in \mathcal{V}$. Note that (i) in Definition \ref{PADICDEF} says that adjoint, if exists,  is unique. An adjointable morphism $U$ is said to be a unitary if $UU^*=U^*U=I_\mathcal{V}$, the identity operator on $\mathcal{V}$.  An adjointable morphism $P$ is said to be projection if $P^2=P^*=P$. An adjointable morphism $T$ is said to be an isometry if $T^*T=I_\mathcal{V}$. An adjointable morphism $T$ is said to be  self-adjoint  if $T^*=T$. We denote the identity operator on $\mathcal{V}$ by $I_\mathcal{V}$. 

Our first result is the indefinite  Halmos dilation. 
\begin{theorem} (\textbf{Indefinite  Halmos dilation})\label{HD}
    	Let $\mathcal{V}$ be a  IIPM over a *-ring of characteristic 2 and 	$T: \mathcal{V} \to \mathcal{V}$ be a self-adjoint morphism. Then the morphism 
    	\begin{align*}
    		U\coloneqq \begin{pmatrix}
    			T & I_\mathcal{V}+T   \\
    			I_\mathcal{V}+T & T  \\
    		\end{pmatrix}
    	\end{align*}
    	is unitary  on 	$\mathcal{V}\oplus \mathcal{V}$. In other words,
    	\begin{align*}
    		T=P_\mathcal{V}U|_\mathcal{V}, \quad 	T^*=P_\mathcal{V}U^*|_\mathcal{V},
    	\end{align*}
    	where $P_\mathcal{V}:\mathcal{V}\oplus \mathcal{V}\ni (x, y) \mapsto x \in \mathcal{V}$.
    \end{theorem}
\begin{proof}
A direct calculation says that 	
\begin{align*}
	V\coloneqq \begin{pmatrix}
		T & I_\mathcal{V}+T   \\
		I_\mathcal{V}+T & T  \\
	\end{pmatrix}
\end{align*}
is the inverse and adjoint of $U$.	
\end{proof}
Our second result is the indefinite Egervary N-dilation.
\begin{theorem} \label{ED}(\textbf{Indefinite Egervary N-dilation})
Let $\mathcal{V}$ be a IIPM over a *-ring of characteristic 2 and $T:\mathcal{V}\to \mathcal{V}$ be a self-adjoint morphism. Let $N$ be a natural number. Then the morphism
\begin{align*}
U\coloneqq \begin{pmatrix}
T & 0& 0 & \cdots &0&0 & I_\mathcal{V}+T   \\
I_\mathcal{V}+T & 0& 0 & \cdots &0&0& T   \\
0&I_\mathcal{V}&0&\cdots &0&0& 0\\
0&0&I_\mathcal{V}&\cdots &0&0 & 0\\
\vdots &\vdots &\vdots & &\vdots & \vdots &\vdots \\
0&0&0&\cdots &0&0 & 0\\
0&0&0&\cdots &I_\mathcal{V}&0 & 0\\
0&0&0&\cdots &0&I_\mathcal{V} & 0\\
\end{pmatrix}_{(N+1)\times (N+1)}
\end{align*}
is unitary on 	$\oplus_{k=1}^{N+1} \mathcal{V}$ and 
\begin{align}\label{FINITEDILATIONEQUATION}
T^k=P_\mathcal{V}U^k|_\mathcal{V},\quad \forall k=1, \dots, N, \quad (T^*)^k=P_\mathcal{V}(U^*)^k|_\mathcal{V},\quad \forall k=1, \dots, N,
\end{align}
where $P_\mathcal{V}:\oplus_{k=1}^{N+1} \mathcal{V} \ni (x_k)_{k=1}^{N+1} \mapsto x_1 \in \mathcal{V}$. 		
\end{theorem}
\begin{proof}
A direct calculation of power of $U$ gives Equation (\ref{FINITEDILATIONEQUATION}). To complete the proof, now we need show that $U$ is unitary. Define 
\begin{align*}
	V\coloneqq \begin{pmatrix}
		T & I_\mathcal{V}+T& 0 & \cdots &0&0 & 0   \\
		0 & 0& I_\mathcal{V} & \cdots &0&0& 0   \\
		0&0&0&\cdots &0&0& 0\\
		0&0&0&\cdots &0&0 & 0\\
		\vdots &\vdots &\vdots & &\vdots & \vdots &\vdots \\
		0&0&0&\cdots &0&I_\mathcal{V} & 0\\
		0&0&0&\cdots &0&0 & I_\mathcal{V}\\
		I_\mathcal{V}+T&T&0&\cdots &0&0 & 0\\
	\end{pmatrix}_{(N+1)\times (N+1)}.
\end{align*}
Then $UV=VU=I_{\oplus_{k=1}^{N+1} \mathcal{V}}$ and  $U^*=V$.
\end{proof}
  Note that the Equation (\ref{FINITEDILATIONEQUATION}) holds only upto $N$ and not for $N+1$ and higher natural numbers. 
 In the following theorem, given a IIPM $\mathcal{V}$,  $\oplus_{n=-\infty}^{\infty} \mathcal{V}$ is the IIPM defined by 
\begin{align*}
\oplus_{n=-\infty}^{\infty} \mathcal{V}\coloneqq \{ \{x_n\}_{n=-\infty}^\infty, x_n \in \mathcal{V}, \forall n \in \mathbb{Z}, x_n\neq 0 \text{ only for finitely many } n'\text{s}\}
\end{align*}
equipped with  inner product 
\begin{align*}
	\langle \{x_n\}_{n=-\infty}^\infty, \{y_n\}_{n=-\infty}^\infty\rangle \coloneqq \sum_{n=-\infty}^{\infty}\langle x_n, y_n \rangle , \quad \forall \{x_n\}_{n=-\infty}^\infty, \{y_n\}_{n=-\infty}^\infty\in \oplus_{n=-\infty}^{\infty} \mathcal{V}.
\end{align*}
Our third result is the indefinite  Sz.-Nagy dilation.
\begin{theorem}\label{ND} (\textbf{Indefinite Sz.-Nagy dilation})
Let $\mathcal{V}$ be a IIPM  over a *-ring of characteristic 2 and 	$T: \mathcal{V} \to \mathcal{V}$ be a self-adjoint morphism.  Let $U\coloneqq(u_{n,m})_{-\infty \leq n,m\leq \infty}$ be the morphism defined on 
$\oplus_{n=-\infty}^{\infty} \mathcal{V}$ given by  the infinite matrix defined as follows:
\begin{align*}
&u_{0,0}\coloneqq T, \quad u_{0,1}\coloneqq I_\mathcal{V}+T, \quad u_{-1, 0}\coloneqq I_\mathcal{V}+T, \quad u_{-1, 1}\coloneqq T, \\
& u_{n,n+1}\coloneqq I_\mathcal{V}, \quad \forall n \in \mathbb{Z}, n\neq 0,1,  \quad u_{n,m}\coloneqq 0 \quad  \text{ otherwise},
\end{align*}
i.e., 
\begin{align*}
U=\begin{pmatrix}
 &\vdots &\vdots & \vdots & \vdots & \vdots &\vdots & \\
 \cdots & I_\mathcal{V}&0&0&0& 0&0&\cdots &\\
\cdots & 0 & I_\mathcal{V} & 0 & 0&  0&0& \cdots & \\
\cdots & 0 & 0& I_\mathcal{V}+T & T& 0&0&\cdots  & \\
\cdots & 0&0&\boxed{T}&I_\mathcal{V}+T& 0&0&\cdots&\\
\cdots & 0&0&0&0& I_\mathcal{V}& 0&\cdots &\\
\cdots & 0&0&0&0& 0&I_\mathcal{V}&\cdots &\\
 & \vdots &\vdots &\vdots &\vdots  &\vdots & \vdots & \\
\end{pmatrix}_{\infty\times \infty}
\end{align*}
where $T$ is in the $(0,0)$  position (which is boxed), is unitary   on 	$\oplus_{n=-\infty}^{\infty} \mathcal{V}$ and 
\begin{align}\label{INFINITEDILATIONEQUATION}
T^n=P_\mathcal{V}U^n|_\mathcal{V},\quad \forall n\in \mathbb{N}, \quad (T^*)^n=P_\mathcal{V}(U^*)^n|_\mathcal{V},\quad \forall n\in \mathbb{N},
\end{align}
where $P_\mathcal{V}:\oplus_{n=-\infty}^{\infty} \mathcal{V}\ni  (x_n)_{n=-\infty}^{\infty} \mapsto x_0 \in \mathcal{V}$.
\end{theorem}
\begin{proof}
We  get Equation (\ref{INFINITEDILATIONEQUATION}) by calculation of powers of $U$. The matrix   $V\coloneqq(v_{n,m})_{-\infty \leq n,m\leq \infty}$ defined by  
\begin{align*}
	&v_{0,0}\coloneqq T, \quad v_{0,-1}\coloneqq I_\mathcal{V}+T, \quad v_{1, 0}\coloneqq I_\mathcal{V}+T, \quad v_{1, -1}\coloneqq T, \\
	& v_{n,n-1}\coloneqq I_\mathcal{V}, \quad \forall n \in \mathbb{Z}, n\neq 0,1,  \quad v_{n,m}\coloneqq 0 \quad  \text{ otherwise},
\end{align*}
i.e., 
\begin{align*}
V=\begin{pmatrix}
&\vdots &\vdots &\vdots & \vdots & \vdots & \vdots & \\
\cdots&I_\mathcal{V} & 0 & 0& 0 & 0&  0& \cdots & \\
\cdots &0& I_\mathcal{V} & 0& 0 & 0&  0& \cdots & \\
\cdots &0& 0 & I_\mathcal{V}+T& \boxed{T} & 0& 0&\cdots  & \\
\cdots &0& 0&T&I_\mathcal{V}+T&0& 0&\cdots&\\
\cdots &0& 0&0&0&I_\mathcal{V}& 0&\cdots &\\
\cdots &0& 0&0&0&0& I_\mathcal{V}&\cdots &\\
&\vdots &\vdots &\vdots &\vdots &\vdots  & \vdots & \\
\end{pmatrix}_{\infty\times \infty}
\end{align*}
where $T$ is in the $(0.0)$  position (which is boxed), satisfies $UV=VU=I_{\oplus_{n=-\infty}^{\infty} \mathcal{V}}$ and  $U^*=V$. 
\end{proof}
We note that explicit sequential form of $U$ is
\begin{align*}
	U(x_n)_{n=-\infty}^{\infty}=(\dots, x_{-2}, x_{-1}, (I_\mathcal{V}+T)x_0+Tx_1, \boxed{Tx_0+(I_\mathcal{V}+T)x_1}, x_2, x_2, \dots)
\end{align*}
where $Tx_0+(I_\mathcal{V}+T)x_1$ is in the $0$  position (which is boxed)
and $U^*$ is 
\begin{align*}
	U^*(x_n)_{n=-\infty}^{\infty}=(\dots, x_{-3}, x_{-2},  \boxed{(I_\mathcal{V}+T)x_{-1}+Tx_0}, Tx_{-1}+(I_\mathcal{V}+T)x_0, x_1, \dots),
\end{align*}
where $(I_\mathcal{V}+T)x_{-1}+Tx_0$ is in the $0$  position (which is boxed).
We next wish to derive indefinite isometric Sz.-Nagy dilation.
\begin{theorem}\label{ISOSZNAGY} (\textbf{Indefinite isometric Sz.-Nagy dilation})
Let $\mathcal{V}$ be a IIPM over a *-ring of characteristic 2 and $T:\mathcal{V}\to \mathcal{V}$ be a self-adjoint morphism.	Let $U\coloneqq(u_{n,m})_{0\leq n,m\leq \infty}$ be the morphism defined on $\oplus_{n=0}^{\infty} \mathcal{V}$ given by  the infinite matrix defined as follows:
\begin{align*}
	u_{0,0}\coloneqq T, \quad u_{2,1}\coloneqq I_\mathcal{V}+T, \quad 
	 u_{n+1, n}\coloneqq I_\mathcal{V}, \quad \forall n \geq 2,  \quad u_{n,m}\coloneqq 0 \quad  \text{ otherwise},
\end{align*}
i.e., 
\begin{align*}
	U=\begin{pmatrix}
		 &\boxed{T}&0&0&0& 0&0&\cdots &\\
		 & I_\mathcal{V}+T & 0 & 0 & 0&  0&0& \cdots & \\
		 &0 & I_\mathcal{V}& 0& 0& 0&0&\cdots  & \\
		 &0&0&I_\mathcal{V}&0& 0&0&\cdots&\\
		 &0&0&0& I_\mathcal{V}&0 & 0&\cdots &\\
		 &0&0&0&0& I_\mathcal{V}&0&\cdots &\\
		&\vdots&\vdots &\vdots &\vdots  &\vdots & \vdots & \\
	\end{pmatrix}_{\infty\times \infty}
\end{align*}
where $T$ is in the $(0,0)$  position (which is boxed), is isometry    on 	$\oplus_{n=0}^{\infty} \mathcal{V}$ and 
\begin{align}
	T^n=P_\mathcal{V}U^n|_\mathcal{V},\quad \forall n\in \mathbb{N}, \quad (T^*)^n=P_\mathcal{V}(U^*)^n|_\mathcal{V},\quad \forall n\in \mathbb{N},
\end{align}
where $P_\mathcal{V}:\oplus_{n=0}^{\infty} \mathcal{V}\ni  (x_n)_{n=0}^{\infty} \mapsto x_0 \in \mathcal{V}$.
\end{theorem}
\begin{proof}
It suffices to note the adjoint of $U$	is 
	\begin{align*}
		U^*=\begin{pmatrix}
			&\boxed{T}&I_\mathcal{V}+T&0&0& 0&0&\cdots &\\
			& 0 & 0 & I_\mathcal{V} & 0&  0&0& \cdots & \\
			&0 & 0& 0& I_\mathcal{V}& 0&0&\cdots  & \\
			&0&0&0&0& I_\mathcal{V}&0&\cdots&\\
			&0&0&0& 0&0 & I_\mathcal{V}&\cdots &\\
			&0&0&0&0& 0&0&\cdots &\\
			&\vdots&\vdots &\vdots &\vdots  &\vdots & \vdots & \\
		\end{pmatrix}_{\infty\times \infty}
	\end{align*}
where $T$ is in the $(0,0)$  position (which is boxed).
\end{proof}
We now formulate following problems.
\begin{problem}
\textbf{\begin{enumerate}[\upshape(i)]
		\item Whether there is an indefinite Ando dilation? If yes, whether  one can dilate commuting three, four, ... commuting self-adjoint morphisms to commuting unitaries?
		\item Whether there is (a kind of) uniqueness of indefinite  Halmos dilation?
		\item Whether there is a indefinite  intertwining-lifting theorem (commutant lifting theorem)?
	\end{enumerate}}
\end{problem}

  \bibliographystyle{plain}
 \bibliography{reference.bib}

\end{document}